# Yoga for Fourier–Mukai partnership

### v2

@@ 4. Fully faithfulness & equivalences
Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ obtained in Proposition˜3.8. Set $\eta\colon 1\to\Phi_{K}\Phi_{K^{\prime}}$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. Then the following are equivalent:
@@
    $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$
@@
    $\eta_{E}$ is an isomorphism for all $E\in D^{-}_{\operatorname{coh}}(Y_{1})$
@@
    $\eta_{E}$ is an isomorphism for all $E\in D^{b}_{\operatorname{coh}}(Y_{1})$
@@
    $\eta_{E}$ is an isomorphism for all $E\in\operatorname{Perf}(Y_{1})$
Additionally, if $K$ is relatively perfect over $Y_{1}$, then these are equivalent to $\Phi_{K}$ is fully faithful on $D^{b}_{\operatorname{coh}}$.
@@
Remark 4.5.
In Proposition˜4.4, a reasonable generalization of the last claim could be to replace $D^{b}_{\operatorname{coh}}$ with $\operatorname{Perf}$ or $D^{-}_{\operatorname{coh}}$.
Theorem 4.6.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. If $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D_{\operatorname{qc}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
First, we check the case of fully faithfulness. Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ on $D_{\operatorname{qc}}$ that is obtained in Proposition˜3.8. Recall that Proposition˜3.8 ensures $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D_{\operatorname{qc}}$. Moreover, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ preserves small coproducts as all integral transforms do. Let $\mathcal{T}$ be the subcategory of $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for which the unit of $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is an isomorphism. By Lemma˜4.3, $\mathcal{T}$ is a triangulated subcategory which is closed under small coproducts. Using Proposition˜4.2, we see that $\mathbf{L}t_{1}^{\ast}\operatorname{Perf}(Y_{1})\subseteq\mathcal{T}$. However, $t_{1}$ is affine, so $\mathbf{L}t_{1}^{\ast}P$ is a compact generator of $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ if $P$ is such for $D_{\operatorname{qc}}(Y_{1})$ (see e.g. [Sta25, Tag 0BQT]). Hence, Proposition˜4.4 ensures that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful on $D_{\operatorname{qc}}$ as desired. Indeed, $\mathcal{T}$ containing $\mathbf{L}t_{1}^{\ast}P$ ensures it contains all the compact objects because
$\operatorname{Perf}(Y_{1}\times_{S}T)=\langle\mathbf{L}t_{1}^{\ast}P\rangle\subseteq\mathcal{T}.$
To see the the inclusion, use that any triangulated subcategory closed under small coproducts is closed under direct summands (e.g. via the ‘Eilenberg swindle’).
Lastly, we check the case of $\Phi_{K}$ being an equivalence. From the first claim, it suffices to show $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective. If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then one can show that $\Phi_{K^{\prime}}$ is also left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. By Proposition˜3.8, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Moreover, as $\Phi_{K^{\prime}}$ is an equivalence on $D_{\operatorname{qc}}$, it must also restrict to an equivalence on $\operatorname{Perf}$. Hence, $K^{\prime}$ must be relatively perfect over $Y_{1}$. Now, by the above $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is fully faithful and by Proposition 4.4 the unit of the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an isomorphism. However, this implies $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective, which completes the proof. ∎
Corollary 4.7.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension and $S$ be semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
First, we show the case of being fully faithful. By Proposition˜4.4, $\Phi_{K}$ restricting to a fully faithful functor on $D^{b}_{\operatorname{coh}}$ is equivalent to $\Phi_{K}$ being fully faithful on $D_{\operatorname{qc}}$ because $K$ is relatively perfect over each $Y_{i}$. So, the desired claim follows from Theorem˜4.6 and a further application of Proposition˜4.4.
Next, we check the case of $\Phi_{K}$ being an equivalence on $D^{b}_{\operatorname{coh}}$. By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies $\Phi_{K}$ is an equivalence on $D_{\operatorname{qc}}$. From Theorem˜4.6, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an equivalence on $D_{\operatorname{qc}}$. However, once again, Remark˜2.7 tells us $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ as desired. ∎
@@ 5. Applications
In this section, we study the consequences of Theorem˜4.6. By ‘genus’ of a curve, we mean ‘arithmetic’(see e.g. [Sta25, Tag 0BY6]).
Proposition 5.1.
Proof.
Let $\overline{k}$ be the algebraic closure of $k$. Suppose $K\in D^{b}_{\operatorname{coh}}(C\times_{k}C^{\prime})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$. Denote by $\pi\colon(C\times_{k}C^{\prime})\times_{k}\operatorname{Spec}(\overline{k})\to C\times_{k}C^{\prime}$ the natural morphism. From Theorem˜4.6, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Then $C\times_{k}\operatorname{Spec}(\overline{k})$ and $C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ are Fourier–Mukai partners over an algebraically closed field of characteristic zero. By [CNS25, Corollary 5.4], we have that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $\operatorname{Perf}(C\times_{k}\operatorname{Spec}(\overline{k}))\to\operatorname{Perf}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Consequently, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ in the following cases:
@@
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
@@
Let $S$ be a quasi-compact separated regular $\mathbb{Q}$-scheme. Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves. Assume $f_{1}$ is locally projective. If $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$, then then genii of the special fibers of $Y_{1}$ coincide with that of $Y_{2}$.
@@
implies that $\mathbb{Q}\subseteq\kappa(s)$ (i.e. $\kappa(s)$ has characteristic zero). Note that, as $S$ is separated, $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g. [Sta25, Tag 01SG]). Denote by $Y_{i}^{s}$ for the base change of $Y_{i}$ along $\operatorname{Spec}(\kappa(s))\to S$. Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral proper curve over $\kappa(s)$. Then, by Theorem˜4.6, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$. Indeed, the hypothesis on $S$ ensures it is Noetherian, and so the closed immersion $\operatorname{Spec}(\kappa(s))\to S$ is locally of finite tor-dimension (see e.g. [GW23, Remark 25.39(2)]). Furthermore, Proposition˜5.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$. Hence, if coupled with [Sta25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$. In other words, the genii of $Y^{s}_{1}$ and $Y_{2}^{s}$ coincide, which completes the proof. ∎
Proposition 5.3.
@@
By Theorem˜4.6, $Y_{1}\times_{k}\operatorname{Spec}(L)$ and $Y_{2}\times_{k}\operatorname{Spec}(L)$ are Fourier–Mukai partners for every finitely generated field extension $L/k$. So, $Y_{1}\times_{k}\operatorname{Spec}(L)$ is regular. Then [CNS25, Corollary 5.4] implies the following string of triangulated equivalences:
@@
Corollary 5.4.
Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a quasi-compact separated regular scheme. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
Proof.
We prove only one direction as the other follows nearly verbatim. Assume $f_{1}$ is smooth. It suffices, by [Sta25, Tag 01V8], to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are smooth over $\kappa(s)$ where $s\in S$ is any point. Fix a point $s\in S$ (which need not be closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. A further application Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Indeed, $S$ being regular implies $\mathcal{O}_{S,s}$ is as well, and so the closed immersion from $\operatorname{Spec(\kappa(p))}$ is of finite tor-dimension (see e.g. [GW23, Remark 25.39(2)]). However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth. Hence, the desired claim follows from Proposition˜5.3. ∎
@@
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms with geometrically integral fibers. Suppose $S$ is quasi-compact, separated, and regular. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is Cohen–Macaulay (resp. Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
Proof.
It suffices to show $f_{2}$ is Cohen–Macaulay (resp. Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse. In fact, we argue in a similar fashion to Corollary˜5.4. Moreover, by [Sta25, Tag 045O] (resp. [Sta25, Tag 0C03]), we only need to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are Cohen–Macaulay (resp. Gorenstein) over an algebraic closure of each $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.6. However, it follows (as fibers are geometrically integral) from [HLS09, Theorem 4.4], that these integral projective schemes over $\overline{\kappa(s)}$ enjoy the property that Cohen–Macaulayness (resp. Gorensteinnes) is a derived invariance. This completes the proof. ∎
Under additional conditions, we can improve Proposition˜5.5 in the Gorenstein case.
Proposition 5.6.
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme, where $Y_{1}$ and $Y_{2}$ are of finite Krull dimension, satisfying the resolution property (see e.g. [Sta25, Tag 0F86]). Assume that $Y_{1}$ and $Y_{2}$ are derived equivalences (e.g. the equivalence need not arise from an integral transform). Then $f_{1}$ is Gorenstein if, and only if, $f_{2}$ satisfies the same condition.

### v3

@@ 4. Fully faithfulness & equivalences
Consider ˜3.1 where $t$ is affine. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. Then, after possibly replacing $\Phi_{K^{\prime}}$ by an isomorphic functor, we have a morphism of adjoint pairs as depicted in the diagram
$D_{\operatorname{qc}}(Y_{1}\times_{S}T)$$D_{\operatorname{qc}}(Y_{1})$$D_{\operatorname{qc}}(Y_{2}\times_{S}T)$$D_{\operatorname{qc}}(Y_{2})$$\scriptstyle{\mathbf{R}(t_{1})_{\ast}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}$$\scriptstyle{\Phi_{K}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}}$$\scriptstyle{\mathbf{R}(t_{2})_{\ast}}$$\scriptstyle{\Phi_{K^{\prime}}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
with left and right comparison transformations given respectively by the isomorphisms
$\alpha^{K}\colon\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}\to\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K},$
$(\alpha^{K^{\prime}})^{-1}\colon\mathbf{R}(t_{1})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\to\Phi_{K^{\prime}}\circ\mathbf{R}(t_{2})^{\ast}.$
Proof.
This argument is essentially dual to that of Proposition˜4.7. ∎
Remark 4.11.
Consider the situation as in Proposition˜4.10. Denote $\varepsilon$ and $\varepsilon^{\prime}$ respectively for the counits of the adjunctions $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$. From Lemma˜4.1, it follows that (after possibly replacing $\Phi_{K^{\prime}}$ by an isomorphic functor) the following diagrams commute:
(4.14) $\Phi_{K}\mathbf{L}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$$\mathbf{R}(t_{2})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$$\Phi_{K}\Phi_{K^{\prime}}\mathbf{R}(t_{2})_{\ast}$$\mathbf{R}(t_{2})_{\ast}$$\scriptstyle{\alpha^{K}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}}}$$\scriptstyle{\mathbf{R}(t_{2})_{\ast}(\varepsilon^{\prime})}$$\scriptstyle{\Phi_{K}(\alpha^{K^{\prime}})}$$\scriptstyle{\varepsilon_{\mathbf{R}(t_{2})_{\ast}}}$
(4.15) $\mathbf{R}(t_{1})_{\ast}$$\Phi_{K^{\prime}}\Phi_{K}\mathbf{R}(t_{1})_{\ast}$$\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$$\Phi_{K^{\prime}}\mathbf{R}(t_{2})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}.$$\scriptstyle{\eta_{\mathbf{R}(t_{1})_{\ast}}}$$\scriptstyle{\mathbf{R}(t_{1})_{\ast}(\eta^{\prime})}$$\scriptstyle{\Phi_{K^{\prime}}(\alpha^{K})}$$\scriptstyle{\alpha^{K^{\prime}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}}$
Proposition 4.12.
Consider ˜3.1 where $t$ is affine. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. By Theorem˜3.6, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$. Hence, Proposition˜3.8 tells us that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$. So, from Corollary˜3.9, we know that $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D^{b}_{\operatorname{coh}}$. Now, let $\eta$ and $\varepsilon$ (resp. $\eta^{\prime}$ and $\varepsilon^{\prime}$) be the unit and counit of the adjunction $\Phi_{K}\dashv\Phi_{K^{\prime}}$ (resp. of $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$). Since $t$ is affine, so are $t_{1},t_{2}$. Therefore, by commutativity of Diagram˜4.15 (resp. by Diagram˜4.14), from [STA25, Tag 08I8] we get that $\eta$ (resp. $\varepsilon$) being an isomorphism implies $\eta^{\prime}$ (resp. $\varepsilon^{\prime}$) is an isomorphism. The result then follows from [STA25, Tag 07RB]. ∎
Theorem 4.13.
Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. Then the following are equivalent:
@@
    $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$
@@
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for any affine morphism $t$
@@
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for every morphism $t\colon\operatorname{Spec}(k)\to S$ from a field
@@
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for every closed point $p\in S$ with associated closed immersion $t\colon\operatorname{Spec}(\kappa(p))\to S$
  5. (5)
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for some affine surjection $t$.
@@ 5. Applications
In this section, we study the consequences of Theorem˜4.13. By ‘genus’ of a curve, we mean ‘arithmetic’(see e.g. [STA25, Tag 0BY6]).
Proof of Proposition˜1.4.
Let $\overline{k}$ be the algebraic closure of $k$. Suppose $K\in D^{b}_{\operatorname{coh}}(C\times_{k}C^{\prime})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$. Denote by $\pi\colon(C\times_{k}C^{\prime})\times_{k}\operatorname{Spec}(\overline{k})\to C\times_{k}C^{\prime}$ the natural morphism. From Theorem˜4.13, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Then $C\times_{k}\operatorname{Spec}(\overline{k})$ and $C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ are Fourier–Mukai partners over an algebraically closed field of characteristic zero. By [CNS25, Corollary 5.4], we have that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $\operatorname{Perf}(C\times_{k}\operatorname{Spec}(\overline{k}))\to\operatorname{Perf}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Consequently, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ in the following cases:
@@
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of genus one and trivial canonical line bundle [LÓP14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [CAT82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
@@
Let $S$ be a Noetherian $\mathbb{Q}$-scheme. Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves. If $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$, then then genii of the special fibers of $Y_{1}$ coincide with that of $Y_{2}$.
@@
implies that $\mathbb{Q}\subseteq\kappa(s)$ (i.e. $\kappa(s)$ has characteristic zero). Note that $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g. [STA25, Tag 01SI]). Denote by $Y_{i}^{s}$ for the base change of $Y_{i}$ along $\operatorname{Spec}(\kappa(s))\to S$. Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral projective curve over $\kappa(s)$ (see e.g. [STA25, Tag 0A26]). Then, by Theorem˜4.13, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$. Furthermore, Proposition˜1.4 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$. Hence, if coupled with [STA25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$. In other words, the genii of $Y^{s}_{1}$ and $Y_{2}^{s}$ coincide, which completes the proof. ∎
Lemma 5.2.
@@
By Theorem˜4.13, $Y_{1}\times_{k}\operatorname{Spec}(L)$ and $Y_{2}\times_{k}\operatorname{Spec}(L)$ are Fourier–Mukai partners for every finitely generated field extension $L/k$. So, $Y_{1}\times_{k}\operatorname{Spec}(L)$ is regular. Then [CNS25, Corollary 5.4] implies the following string of triangulated equivalences:
@@
Proof of Proposition˜1.5.
We prove only one direction as the other follows nearly verbatim. Assume $f_{1}$ is smooth. It suffices, by [STA25, Tag 01V8], to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are smooth over $\kappa(s)$ where $s\in S$ is any point. Fix a point $s\in S$ (which need not be closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [STA25, Tag 01SI]). Hence, Theorem˜4.13 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. A further application Theorem˜4.13 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth. Hence, the desired claim follows from Lemma˜5.2. ∎
Proof of Proposition˜1.6.
It suffices to show $f_{2}$ is Cohen–Macaulay (resp. Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse. In fact, we argue in a similar fashion to Proposition˜1.5. Moreover, by [STA25, Tag 045O] (resp. [STA25, Tag 0C03]), we only need to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are Cohen–Macaulay (resp. Gorenstein) over an algebraic closure of each $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [STA25, Tag 01SG]). Hence, Theorem˜4.13 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜4.13 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.13. However, it follows (as fibers are geometrically integral) from [HLS09, Theorem 4.4], that these integral projective schemes over $\overline{\kappa(s)}$ enjoy the property that Cohen–Macaulayness (resp. Gorensteinnes) is a derived invariance. This completes the proof. ∎
@@
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme, where $Y_{1}$ and $Y_{2}$ are of finite Krull dimension, satisfying the resolution property (see e.g. [STA25, Tag 0F86]). Assume that $Y_{1}$ and $Y_{2}$ are derived equivalent (e.g. the equivalence need not arise from an integral transform). Then $f_{1}$ is Gorenstein if, and only if, $f_{2}$ satisfies the same condition.